i am unable to solve circuit problems of resistance, capacitance, inductance and circuits containing all the three.... please give me tips or   steps to solve the circuits........please help me.....

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Section 1. Series Circuit

  Check!
Series Circuit

(Series Circuit)

The total voltage is the sum of the voltage on each component.
eq 1: V0 = V1+ V2 + V3 +...+ Vn
(In this case, VT = V1+ V2)

The total resistance is equal to the sum of the resistance on each component.
eq 2: R0 = R1 + R2 + R3 +...+ Rn
(In this case, RT = R1 + R2)

The total current is equal in every component.
eq 3: I0 = I1 = I2= I3= I4 =...= In
(In this case, IT = I1 = I2)